At 58.0 ∘C , what is the maximum value of the reaction quotient, Q, needed to produce a non-negative Ecell value for the reaction SO42−(aq)+4H+(aq)+2Br−(aq)⇌Br2(l)+SO2(g)+2H2O(l) In other words, what is Q when Ecell=0 at this temperature?

I found that Ecell = 0.87 V and n= 2

−nF(Ecell−E∘cell) / 2.303RT =
2×96485 (0+0.87)/2.303×8.314×331.0

log q = -26.49

I need Q without log
Answer #1

According to the nernest equation,

Ecell = E0cell - (2.303 R*T/(n*F))*logQ

Where Q is the reaction quotient

F is Faraday constant

n is no. of electrons transferred

T is temperature in Kelvin

R is the universal gas constant

Now, as it is given in question,

Ecell = 0

E0cell = 0.87V

n = 2

T = 58°C = 273 + 58 = 331 K

R = 8.314 J/mol K

F = 96485 C/ mol

Substituting values in equation,

0 = 0.87 - (2.303*8.314*331/(2*96485))*log Q

0 = 0.87 - 0.03284 * log Q

Hence , log Q = 0.87 / 0.03284 = 26.49

Q = antilog (Q) = antilog (26.49) = 3.09 * 10-26 (answer )

I hope you will find the answer useful. For any further clarification please ask in comments.